OpenStudy (anonymous):

HELp PLEASE!! Multiply an simplify if possible... sqrt(7x) (sqrt(x) -7 sqrt(7)) I got 336x

OpenStudy (anonymous):

\[\sqrt{7} (\sqrt{x}-7\sqrt{7})\]

OpenStudy (mathmate):

Would it not be \(\sqrt{7x} (\sqrt{x}-7\sqrt{7})\) ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@mathmate

OpenStudy (solomonzelman):

wait, so the problem is: \(\large\color{black}{\sqrt{7x} \times(\sqrt{x} -7\sqrt{7}) }\)

OpenStudy (solomonzelman):

\(\large\color{black}{\sqrt{7x} \times\sqrt{x}=? }\) \(\large\color{black}{\sqrt{7x} \times7\sqrt{7}=? }\)

OpenStudy (anonymous):

yes that's the problem.

OpenStudy (anonymous):

\[\sqrt{7x^2}\]?

OpenStudy (solomonzelman):

for the first one you are correct.

OpenStudy (solomonzelman):

so, can you simplify the first one?

OpenStudy (anonymous):

would it be 7x

OpenStudy (solomonzelman):

not exactly.

OpenStudy (anonymous):

ok then I am not sure

OpenStudy (solomonzelman):

\(\large\color{black}{\sqrt{7x^2}=\sqrt{7}\times\sqrt{x^2} =? }\)

OpenStudy (anonymous):

I have to list my steps for this question

OpenStudy (anonymous):

oh uhm no idea

OpenStudy (solomonzelman):

we will do this when we overview

OpenStudy (solomonzelman):

\(\large\color{black}{\sqrt{x^2} =x }\) So, \(\large\color{black}{\sqrt{7}\times\sqrt{x^2} =? }\)

OpenStudy (solomonzelman):

wouldn't it just be \(\large\color{black}{\sqrt{7}~x }\) ?

OpenStudy (solomonzelman):

(or, \(\large\color{black}{x\sqrt{7} }\), same thing )

OpenStudy (anonymous):

Ok so its x\[\sqrt{7}\]

OpenStudy (solomonzelman):

yes. \(\large\color{black}{x\sqrt{7} }\)

OpenStudy (solomonzelman):

and \(\large\color{black}{\sqrt{7x}\times7\sqrt{7} =? }\)

OpenStudy (anonymous):

no idea

OpenStudy (anonymous):

it went away

OpenStudy (solomonzelman):

I deleted it, because it had an error. \(\large\color{black}{7\sqrt{7}\times\sqrt{7x} }\) I am braking it into, \(\large\color{black}{\sqrt{7}\times\sqrt{7} \times \sqrt{x}\times7 }\)

OpenStudy (solomonzelman):

can you tell me, what \(\large\color{black}{\sqrt{7}\times\sqrt{7} \times \sqrt{x}\times7 }\) will be?

OpenStudy (solomonzelman):

Rule: (for any real or non-real number \(\large\color{black}{c }\) ) \(\large\color{red}{\sqrt{c}\times\sqrt{c} =c }\)

OpenStudy (solomonzelman):

so, what will happen to the \(\large\color{blue}{\sqrt{7}\times\sqrt{7} }\) part of the \(\large\color{black}{\sqrt{7}\times\sqrt{7} \times \sqrt{x}\times7 }\) ?

OpenStudy (anonymous):

7x?

OpenStudy (solomonzelman):

that's the whole thing, \(\large\color{black}{\sqrt{7}\times\sqrt{7} \times \sqrt{x}\times7 }\) ?

OpenStudy (anonymous):

like the answer?

OpenStudy (solomonzelman):

it's not the answer to your (entire) problem.

OpenStudy (anonymous):

ok based on the c thing you said above

OpenStudy (anonymous):

7x^2?

OpenStudy (solomonzelman):

based on you have: \(\large\color{black}{\sqrt{7}\times\sqrt{7} \times \sqrt{x}\times7 }\) \(\large\color{black}{~~~~~~~~7~~~~~~\times \sqrt{x}\times7 }\)

OpenStudy (anonymous):

ok....49x

OpenStudy (solomonzelman):

and \(\large\color{black}{7 \times7=49 }\). So, \(\large\color{black}{49\sqrt{x} }\)

OpenStudy (solomonzelman):

you were close

OpenStudy (anonymous):

oh?

OpenStudy (anonymous):

Yay I almost got something!

OpenStudy (solomonzelman):

you are going well.... I love disconnecting in a middle of a reply, makes me feel excited

OpenStudy (anonymous):

Oh no!

OpenStudy (solomonzelman):

We had: \(\large\color{black}{\sqrt{7x}~(\sqrt{x}~-~7\sqrt{7}) }\) We said: ~ \(\large\color{black}{\sqrt{7x}~\times~\sqrt{x}=x\sqrt{7} }\) ~ \(\large\color{black}{\sqrt{7x}~\times~7\sqrt{7}=49\sqrt{x} }\)

OpenStudy (solomonzelman):

can you tell me the answer to your prblem?

OpenStudy (anonymous):

\[x \sqrt{7} \times 49\sqrt{x}\]

OpenStudy (solomonzelman):

you have a minus, in the parenthesis, so, shouldn't it be, \(\large\color{blue}{x\sqrt{7}- 49\sqrt{x} }\) ?

OpenStudy (anonymous):

ohh! I see

OpenStudy (solomonzelman):

Okay, then you want to show your work, right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

this is what i got

OpenStudy (anonymous):

oh poo it wont let me post

OpenStudy (solomonzelman):

if the post button is hiding, click Tab+enter

OpenStudy (anonymous):

there we go

OpenStudy (solomonzelman):

you are correct.

OpenStudy (anonymous):

OK sweet!!!

OpenStudy (anonymous):

I have 2 left. Can you help with those as well?

OpenStudy (solomonzelman):

what I would do though, is in 2nd and 3rd row, explain that you are solving for each of the products, to the combine the result of what you get after explanding.

OpenStudy (solomonzelman):

*expanding

OpenStudy (anonymous):

ok :)

OpenStudy (solomonzelman):

will see how it goes for me...